OpenStudy (anonymous):

Use the laws of exponents to simply! 9.1^-5/3 --------- 9.1^6/7 The 5/3 the whole thing is negative not just the 5

OpenStudy (anonymous):

hmmm tricky....

myininaya (myininaya):

9.1^(-5/3-6/7)

myininaya (myininaya):

9.1^(-53/21)

myininaya (myininaya):

\[\frac{1}{9.1^\frac{53}{21}}\]

myininaya (myininaya):

since 9.1=91/10 \[\frac{10^\frac{53}{21}}{91^\frac{53}{21}}\]

OpenStudy (anonymous):

does that say 53/21

OpenStudy (anonymous):

i 1/9.1/83/21

myininaya (myininaya):

yes it says 53/21